40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Mar 2 23:37:36 2012, tara, DailyProgress, BEAT, beat is back 
    Reply  Sat Mar 3 14:53:51 2012, Zach, DailyProgress, BEAT, beat is back disco.png
       Reply  Sat Mar 3 17:50:45 2012, tara, DailyProgress, BEAT, beat is back 
          Reply  Sat Mar 3 18:07:15 2012, Zach, DailyProgress, BEAT, beat is back 
    Reply  Sun Mar 4 00:45:26 2012, frank, tara, DailyProgress, BEAT, beat is back beat_2012_02_03.png
       Reply  Mon Mar 5 16:29:50 2012, frank, tara, DailyProgress, BEAT, beat is back relax.JPG
       Reply  Tue Mar 6 20:48:47 2012, frank, tara, DailyProgress, BEAT, beat is back beat_2012_02_22.png
Message ID: 858     Entry time: Fri Mar 2 23:37:36 2012     Reply to this: 860   863
Author: tara 
Type: DailyProgress 
Category: BEAT 
Subject: beat is back 

 Beat signal is back now, but we have not measured the spectrum yet since the temperature is still drifting fast (10kHz/ min). We plan to measure it tomorrow afternoon.

    The optics for RCAV are aligned. The visibility is ~95% without adjusting the lenses for mode-matching. I do not redo the mode match yet because we probably have to move the chamber for installing the air springs very soon. For ACAV, the beam is aligned, with visibility only ~ 80%. I also adjusted the cable length for PDH locking so that the error signal looks symmetric (Frank added the 4-ch splitter for demodulating RFAM, so the phase shift changed a bit.  All reflected beams  are properly blocked with razor blade blocks.

    The beat frequency is ~ 188MHz instead of the nominal value of 160 MHz because the temperature is not settled at the set point yet. It takes longer than before because of the copper shields around the cavities.  We expect it to be stable by tomorrow afternoon.

ELOG V3.1.3-